A Square flooring tile has a circular design imprinted on it. If the tile is 4m long on each side, what is the area of the tile not covereed by the design?

Answer by mananth(16949)

Area of tile - Area of design = area of tile not covered by design
Area of tile = 4*4=16m^2
Area of circular design= pi*r^2
Area = pi*2*2
area = 12.56 m^2
Balance area = 16-12.56=3.44 m^2